OpenStudy (anonymous):

Solve |2y - 2| = y. Answer {2/3, 2} {2, 3} {2, 5} {1/3, 2}

Parth (parthkohli):

\(\Rightarrow 2y - 2 = \pm y \) This is a good start

jimthompson5910 (jim_thompson5910):

Or you can think of it as |2y - 2| = y breaking up into 2y - 2 = y or 2y - 2 = -y

OpenStudy (anonymous):

would it be (1/3, 2)

jimthompson5910 (jim_thompson5910):

no, part of that is correct though

jimthompson5910 (jim_thompson5910):

ie, one of those numbers is correct

OpenStudy (anonymous):

the 2 is correct.

Parth (parthkohli):

Yes, 2 is correct.

Parth (parthkohli):

Now start with the second equation.

OpenStudy (anonymous):

Okay so now im getting (2/3, 2)

Parth (parthkohli):

Yep!